OpenStudy (anonymous):

Calculate the highest frequency X-rays produced by 8.00 · 104 eV electrons.

OpenStudy (shamim):

i think u hv to use equation E=hf E=104ev=104*1.6*10^-19J h=6.63*10^-34J.S now just find frequency of electron

OpenStudy (shamim):

m i right?

OpenStudy (anonymous):

Okay, I thought thats what I needed to use.

OpenStudy (anonymous):

Although I'm no more sure than you are. ;)

OpenStudy (shamim):

ok